Income statement: In an accounting conference, discussion turned to the possibility of preparing financial statements from a few key accounts together with financial or cost ratios The assistant controller of a participating firm provided the following data pretax income for the year $1,200,000; pretax income rate on sales 10% gross profit rate 40% rate of marketing expenses to sales 15%; 5% bonds payable represent 37.5% of the total liabilities of $ 2,000,000

Required: An income statement for the year based on the above information
Answer #1

Answer:

Pretax Income (EBT) = $1,200,000

Pretax income rate on sales is 10%.

Hence sales = Pretax Income / Pretax income rate = $1,200,000 / 10% = $12,000,000

Gross profit rate 40% to sales

Hence, Gross profit = Sales * Gross profit rate = $12,000,000 * 40% = $4,800,000

Cost of goods sold = Sales - Gross profit = $12,000,000 - $4,800,000 = $7,200,000

Rate of marketing expenses to sales 15%

Hence Marketing expense = Sales * Rate of marketing expenses to sales =$12,000,000 * 15% = $1,800,000

5% bonds payable represent 37.5% of the total liabilities of $ 2,000,000

Hence, bonds payable = $2,000,000 * 37.5% = $750,000

Interest expense = Bonds payable * Interest rate = $750,000 * 5% = $37,500

EBIT = EBT + Interest expense = $1,200,000 + $37,500 = $1,237,500

Other operating expense = Gross profit - Marketing expense - EBIT = $4,800,000 - $1,800,000 - $1,237,500 = $1,762,500

Income statement is as follows:

Statement of Income $12,000,000 $7,200,000 $4,800,000 Cost of Goods Sold Gross Profit Other Operating expenses $1,762,500 Marketing expense EBIT51,800,000 Interest expense Pretax Income (EBT) $37,500 1,200.000